OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1/4x + 2. y = 4x - 1 y = 4x - 3 y = 1/4x - 3 y = 1/4x - 1

OpenStudy (dls):

\[\Huge y=mx+c\] where m is the slope,compare with y = 1/4x + 2 The co-efficient of x is the slope. Now use standard equation, \[\Huge y-y_1=m(x-x_1)\]

OpenStudy (dls):

(x1:-4, y1: -2)

terenzreignz (terenzreignz):

Step by step guys... Slopes, @stephanie12212 When two lines are parallel, what's the relation of their slopes?

OpenStudy (anonymous):

so @DLS how do I solve that after I put those in the equation?

OpenStudy (dls):

that is your answer..

OpenStudy (dls):

 Parallel lines have same inclination/slope.

OpenStudy (dls):

 try solving now